OpenStudy (anonymous):

factor completely (5x^4-44x^2+63)

OpenStudy (anonymous):

X= +-3/sqrt(5), x=+-sqrt(7)

OpenStudy (ranga):

Make the substitution t = x^2 The problem will now become a quadratic expression in t. See if you can factor it. Later put back x^2 in pace of t and see if it can be factored further.

OpenStudy (anonymous):

Yup sub in something for x^2

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

im confused

OpenStudy (ranga):

Let t = x^2 5x^4-44x^2+63 becomes 5t^2 - 44t + 63 Now this is a quadratic equation in t. See if it can be factored.

OpenStudy (anonymous):

u multiply 5 n 63 right

OpenStudy (ranga):

You can try this: 5t^2 - 44t + 63 = 5t^2 - 35t - 9t + 63 Now try factoring it.

OpenStudy (anonymous):

wat do u do after that thats the point im at n dont understand it

OpenStudy (ranga):

5t^2 - 44t + 63 = 5t^2 - 35t - 9t + 63 = 5t(t - 7) - 9(t - 7) = (5t - 9)(t - 7) Now put back t = x^2 (5x^2 - 9)(x^2 - 7) Do you know if they allow you to use radicals in factors? If allowed we can factor this further.

OpenStudy (anonymous):

wats radicals

OpenStudy (ranga):

symbols such as the square root sign, cube root, etc.

OpenStudy (anonymous):

idk wat that is

OpenStudy (ranga):

If you look in the chapter that asked this question and you see some examples where they have factored and radical symbols appear in those factors then we can do the same here and factor further.

OpenStudy (anonymous):

o yea i think u do have to do that

OpenStudy (ranga):

Okay in that case let me introduce radical symbols: There is an identity for difference of two squares: a^2 - b^2 = (a + b)(a - b) So we need to put each factor in \[\Large (5x^2 - 9)(x^2 - 7) = ((\sqrt{5}x)^2 - 3^2)(x^2 - \sqrt{7}^2)\]as a difference of two squares as shown:

OpenStudy (anonymous):

huh

OpenStudy (ranga):

And the answer will be: \[\Large (\sqrt{5}x + 3)(\sqrt{5}x - 3)(x + \sqrt{7})(x - \sqrt{7})\]

OpenStudy (ranga):

okay, we have this identity called difference of two squares: a^2 - b^2 = (a + b)(a - b) So if we can put something that looks like the expression on the left then we can factor it the way it appears on the right. Just to give you an example: If we see (x^2 - 4) we can write is as: (x^2 - 2^2) Now this is a difference of two squares. Compare this to a^2 - b^2 and you can see x is a and 2 is b. So it can be factored as: (a + b)(a - b) or (x + 2)(x - 2). So (x^2 - 2^2) = (x + 2)(x - 2) Follow so far?

OpenStudy (anonymous):

yea

OpenStudy (ranga):

So we have two factors: (5x^2 - 9) and (x^2 - 7) Let us take the first factor (5x^2 - 9). I want to make this as a difference of two squares. So you take the square root of the first term and then square it. You take the square root of the second term and then square it. Square root and squares will cancel out and so we won't be changing the expression. It is like writing 16 as square of (square root of 16). The square and square root will cancel out leaving with just 16 as before. So 5x^2 can be written as square of (square root of 5x^2). square root of 5x^2 = square root of 5 * square root of x^2 = \[\sqrt{5} * x\]

OpenStudy (ranga):

9 can be written as square of (square root of 9) = square of (3) Therefore, \[(5x^2 - 9) = (\sqrt{5}x)^2 - 3^2 = (\sqrt{5}x + 3)(\sqrt{5}x - 3)\]

OpenStudy (anonymous):

ok

OpenStudy (ranga):

\[(x^2 - 7) = x^2 - (\sqrt{7})^2 = (x + \sqrt{7})(x - \sqrt{7})\]

OpenStudy (anonymous):

ok

OpenStudy (ranga):

\[\Large (\sqrt{5}x + 3)(\sqrt{5}x - 3)(x + \sqrt{7})(x - \sqrt{7})\]is the result when we factor the original problem completely.

OpenStudy (anonymous):

ok thanks

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

this is easy after u gt the hang of it